# Empirical Validation of the Independent Chip Model

### v1

@@ I Introduction
In this paper, we make two contributions. First, we release a dataset [^1] Our dataset is available on Zenodo: https://zenodo.org/records/15556774. of poker tournaments scraped from WSOP.com (before their recent website redesign), yielding results of 9,958 events, payouts of 5,018 events, and end-of-day chip count summaries of 1,809 events, and PokerNews, providing end-of-day chip count summaries of 3,610 events and payouts of 4,024 events. Our new dataset may allow future researchers to unlock extra insights related to poker tournaments. Second, from chip counts filtered from these two sources (2,500 relevant end-of-day chip count snapshots and results, totaling 33,478 players), we perform a large-scale empirical validation of the ICM. In our first experiment, we compare the ICM’s performance with a baseline we propose, finding that the ICM is indeed more accurate. Next, in our second experiment, we attempt to settle the inconsistencies between the observations of Scott et al. [14] and Henke [16]. We find that the ICM tends to give overly pessimistic expectations for players with larger stacks while being overly optimistic in its predictions of short stacks’ performances, in line with Henke’s observations.

### v2

@@ I Introduction
In this paper, we make two contributions. First, we release a dataset of poker tournaments scraped from WSOP.com (before their recent website redesign), yielding results of 9,958 events, payouts of 5,018 events, and end-of-day chip count summaries of 1,809 events, and PokerNews, providing end-of-day chip count summaries of 3,610 events and payouts of 4,024 events. Our new dataset may allow future researchers to unlock extra insights related to poker tournaments. Second, from chip counts filtered from these two sources (2,500 relevant end-of-day chip count snapshots and results, totaling 33,478 players), we perform a large-scale empirical validation of the ICM. In our first experiment, we compare the ICM’s performance with a baseline we propose, finding that the ICM is indeed more accurate. Next, in our second experiment, we attempt to settle the inconsistencies between the observations of Scott et al. [14] and Henke [16]. We find that the ICM tends to give overly pessimistic expectations for players with larger stacks while being overly optimistic in its predictions of short stacks’ performances, in line with Henke’s observations.